With an 18km/h tail wind a crow flies to a pond in 18 minutes. Find the crow's speed ( still in the air) if the return trip, with the head wind requires 46 minutes.
Answer by Alan3354(69443)

With an 18km/h tail wind a crow flies to a pond in 18 minutes. Find the crow's speed ( still in the air) if the return trip, with the head wind requires 46 minutes.
------------
C = crow's airspeed
W = wind speed
Going downwind, the ground speed is the sum of the 2. Going upwind, it's the difference.
The time is the distance, s, divided by the ground speed, or
t = s/v (s is distance, v = ground speed)
Going with the wind:
18 = s/(C+W) = s/(C+18)
Going against the wind:
46 = s/(C-W) = s/(C-18)
-----------------
s = 18C + 324
s = 46C - 468
Subtract
0 = -28C + 792
28C = 792
C = 792/28 = 198/7 km/hr
----------------
The distance can be solved for also, but if you have this on a test and time is limited, don't waste time doing that. Give them what is asked for and go on to the next problem.